OpenStudy (rainbow_rocks03):

What is the product of 4 2/5 and 2 1/7?

OpenStudy (alexandervonhumboldt2):

first convert into improper fractions: \[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\] \[4\frac{ 2 }{ 5 }=\frac{ 4*5+2 }{ 5 }=\frac{ 22 }{ 5 }\] \[2\frac{ 1 }{ 7 }=\frac{ 2*7+1 }{ 7 }=\frac{ 15 }{ 7 }\] Now multply: \[\frac{ 22 }{ 5 }*\frac{ 15 }{ 7 }=\frac{ 22*15 }{ 5*7 }=simplify\]

OpenStudy (rainbow_rocks03):

22*15 = 330 & 5*7 = 35

OpenStudy (rainbow_rocks03):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

\[\frac{ 330 }{ 35 }=9\frac{ 15 }{ 35 }=9\frac{ 3 }{ 7 }\]

OpenStudy (rainbow_rocks03):

TY!!!